OpenStudy (phebe):

SOLVE the equation using the Zero Product property. (2x-4)(2x-1)=0 A-2 and -1/2 B-2 and 1/2 C- -2, and 2 D- -2 and 1/2

OpenStudy (phebe):

@mjoxprm360

OpenStudy (anonymous):

can you solve \[2x-4=0\]?

OpenStudy (phebe):

yes

OpenStudy (anonymous):

Set each factor to zero and solve for x... 2x - 4 = 0 and 2x - 1 = 0 x = {4/2, ½} x = {2,½}

OpenStudy (phebe):

ok its B

OpenStudy (anonymous):

:)

OpenStudy (phebe):

XD